# The 2010 Census Confidentiality Protections Failed, Here’s How and Why

### v1

@@ 4.6. Solvar Implementation Details
The differences from the reconstruction in equation (4) are that equality constraints for tract-level tables become inequality constraints, and the other blocks in tract $t$ are ignored when computing a reconstruction for block $b^{\prime}$. Specifically, only optimization variables for block $b^{\prime}$ are used when forming the tract-level constraints, hence the left-hand side of the tract-level constraint is an underestimate of the counts in the tract-level tables.
@@
Note that the tract-level tables in the solvar equation (10) appear in inequality constraints. Variables for records in other blocks $b\neq b^{\prime}$ within the same tract $t$ do not appear in this optimization problem for block $b^{\prime}$.
@@
We note that the IP for the initial solution for $\text{rHDF}^{S0}$ is a strict relaxation of the constraints used for the reconstruction $\text{rHDF}_{b,t}$ solving equation (4). This implies that the feasible region for $\text{rHDF}^{S0}$ is a superset of the feasible region for $\text{rHDF}_{b,t}$. Similarly, the solvar IP in equation (10) also searches for $\text{rHDF}^{\prime}$ from a strictly larger solution space. This property implies that $\text{max}(2cumsolvar(\text{rHDF}^{S0},\text{B}_{t}),100)$ upper-bounds the solvar value for tract $t$ that an attacker would get using the stronger tract-level constraints. In other words, $\text{max}(2cumsolvar(\text{rHDF}^{S0}$, $\text{B}_{t})$, 100)100)100 ) upper-bounds the percentage of records by which any two solutions to $\text{rHDF}_{b,t}$ can differ because first, the union $\cup_{b\in\text{B}_{t}}\text{rHDF}^{S0}(b)$ is a solution to a problem with a strictly larger solution space than $\text{rHDF}_{b,t}$, and second, $2\sum_{b\in\text{B}_{t}}solvar(\text{rHDF}^{S0},b)$ upper-bounds the percentage of records by which any two solutions to this problem can disagree.
@@
The problem relaxations we performed probably made our upper bound on solution variability much looser. Hence, strengthened (and even more concerning) solution variability bounds may be produced in the future. However, since summing the block-level solution to equation (10) and multiplying by 2 is also an upper bound on the exact solution variability for the tract-level problem, the large number of blocks with 0 block-level solution variability will still contribute 0 to the tract-level solution variability in the $agebin$ schema. Thus, the variability in these blocks is already bounded by 0 and cannot get tighter. Tracts with 0 tract-level solution variability found in solving the optimization problem (10) must also have 0 solution variability in any stronger tract-level IP.
@@
We found that solution variability could be readily computed using the optimization in equation (10) for all 6,207,027 blocks with positive population in the 2010 Census tabulations. This demonstrates that attackers have a public, computationally feasible method for independently identifying blocks for which aggregation into tables introduces no additional uncertainty about the underlying microdata beyond the SDL measures that were applied at the record level to generate the HDF. No access to confidential data is required to perform these solution variability calculations. Furthermore, population uniques found in blocks with 0 solution variability are provably population uniques in the HDF, again without using any confidential data. And population uniques using the $agebin$ feature are provably population uniques on the exact age feature. Finally, incorporating additional SF1 tables to the ones shown in Table 2 can only decrease solution variability. Since we used only 18 of the 177 census block-level tables and 16 of the 84 census tract-level tables, our solutions to equation (10) are also upper bounds on the solution variability of reconstructions that use additional tables.
@@ 5. Matching and Reidentification Methodology
For either $\text{rHDF}_{b,t}$ or $\text{rHDF}_{b}$, our first-order assessment of the quality of the reconstructed microdata matches these reconstructions directly to both the HDF and CEF. This is not a record linkage assessment. It compares the reconstructed microdata directly to the source microdata for the contingency tables in SF1, namely HDF, and to the source microdata for HDF, namely CEF. We label this step the agreement match because it provides measures of agreement in record-level feature values between $\text{rHDF}_{b,t}$ or $\text{rHDF}_{b}$ and the confidential HDF and CEF. After the agreement match, we do record-linkage based reidentifcation matches to see how accurately an attacker can infer race and ethnicity from the reconstructed microdata. Thus, we link $\text{rHDF}_{b,t}$ and $\text{rHDF}_{b}$ separately to both commercial data and a specially constructed extract from the CEF called $\text{CEF}_{atkr}$ that includes linking variables (quasi-identifiers) from the schema in equation (2) and the person identifier pik but no other variables—specifically, not race or ethnicity. These feature sets are shown in the rows COMCRL and $\text{CEF}_{atkr}$ on Table 1.
@@
By record-linkage of the quasi-identifiers {block, sex, (age or agebin)} in reconstructed microdata to attacker databases that include names (feature pik), we create putative reidentifications. To enhance the attacker’s database, we attach the data on {race, ethnicity} found on $\text{rHDF}_{b,t}$ or $\text{rHDF}_{b}$ to the {pik, block, sex, (age or agebin)} information in the attacker’s database—either COMCRL or $\text{CEF}_{atkr}$. The feature sets for the records in these putative reidentification databases are shown in the rows labeled Putative $\text{rHDF}_{b,t}$ and Putative $\text{rHDF}_{b}$ in Table 1. Finally, to evaluate the accuracy of the putative reidentifications, and classify confirmed reidentifications, we match putative reidentifications to the full CEF, linking on {pik, block, sex, (age or agebin)} and comparing the {race, ethnicity} inferred from the reconstructed microdata (attached to a putatively reidentified person) to that person’s actual census responses in the CEF. We label the reidentification confirmed, when {pik, block, sex, (age or agebin), race, ethnicity} all match in either the schema in equation (2) for the feature age or the schema in Table 5 for the feature agebin. The feature sets for these confirmed reidentifications are shown in the rows labeled Confirmed $\text{rHDF}_{b,t}$ and Confirmed $\text{rHDF}_{b}$ in Table 1. Finally, we use the ratio of confirmed to putative identifications as the measure of precision or accuracy for the attack.
@@
To make the meaning of this exercise as clear as possible, we state the attacker model we are simulating concisely here. The attacker is an entity external to the Census Bureau with no access to the confidential data contained in HDF or CEF. The attacker has access to all published 2010 Census data—every table in SF1, in particular. The attacker selects a subset (possibly the universe) of these tables and performs record-level reconstruction, possibly using the algorithms in this paper. The attacker also has access to an external database that contains name and address (or some other personal identifying information sufficient to tag a unique person like Social Security Number) and quasi-identifiers—features that match some of the features in SF1—specifically, census block (geocoded from address), sex, and age. Notice that we have explicitly used the same feature set definitions for the attacker’s database as we used in our reconstruction. This is an essential characteristic of a record-linkage attack—the attacker knows enough about the definitions of the feature set for SF1 to construct quasi-identifiers with the same schemas. The attacker matches the external data to the reconstructed record-level data from SF1 using the matching variables (quasi-identifiers) {block, sex, age} and adds the other variables reconstructed from SF1 to the external database. These added variables can include any feature tabulated in SF1, including those tabulated for housing units and householders because those data can be reconstructed as part of the record for “person 1” or “householder,” the individual who completed the census responses for the persons living in the housing unit. In our experiments the extra variables are {race, ethnicity}. [^19] Implementing additional features in the reconstructed data is outside the scope of this paper; however, it is straightforward to add the feature “relationship to householder.” We sketch the details here. See the documentation for Table P29 Household Type by Relationship [sf1:2010:tech]. Notice that the universe is “Total Population,” which implies that the marginal total in each of these block-level tables is the block population in Table P1. Create binary variables for the 20 mutually exclusive and exhaustive relationships defined in Table P29: family male householder, family female householder, spouse, biological child, adopted child, stepchild, grandchild, brother or sister, parent, parent-in-law, son-in-law or daughter-in-law, other relatives, nonrelatives, nonfamily male householder living alone, nonfamily male householder not living alone, nonfamily female householder living alone, nonfamily female householder not living alone, nonfamily nonrelatives, group quarters institutionalized person, and group quarters noninstitutionalized person. Add the statistics from Table P29 to the set in Table 2. Add the constraints implied by these tables to the IP in equations (4) and (10). To further reduce solution variability, if any, in the relationship variables and {race, ethnicity}, add Tables P29A–P29I Household Type by Relationship for major {race, ethnicity} categories. To further reduce the solution variability, if any, in the household population universe, add Table P30. To further reduce the solution variability, if any, for children add Tables P31 and P32. To further reduce the solution variability, if any, for persons ages 65 and older, add Table P34. It is not necessary to add the full variable “relationship to householder” to further reduce solution variability in the features {race, ethnicity}. Using only “lives in household” (all relationships except the last two) and “lives in group quarters” (the last two relationships), use Tables P16A–P16I Population in Households by Age for major {race, ethnicity} categories. Once the relationship “householder” has been reconstructed (all categories that contain the word householder), solution variability for the householder’s {race, ethnicity} can be further reduced by adding Tables P18A–P18I Household Type for major {race, ethnicity} categories. Many tract-level SF1 tables also contain information that can further reduce remaining solution variability.
@@
At this point the attacker is using the information from what we call putative reidentifications as record-level variables on the external database. The attacker may not care about the quality of these putative reidentifications; however, because the solution variability in the schema of our attack is 0 for 70% of all census blocks using just the 34 tables in Table 2, the attacker may be satisfied with the 97 million persons for whom they know with certainty they have exactly the information on the confidential internal database. Alternatively, the attacker can keep adding tables from SF1 until solution variability is 0 for as large a portion of the target population as desired. For all records for which solution variability is 0, the attacker knows with certainty that the data added from SF1 for those persons match the confidential source records exactly. That is, the attacker knows that they have an image of the confidential Hundred-percent Detail File (HDF) that is completely accurate in the schema of the feature set used for SF1 and the attack. Traditional SDL did not anticipate the possibility of reconstructing accurate microdata from published tables. That is the new feature of our experiments. However, the reidentification attacks we simulate using the record-level reconstructed data exactly match the textbook descriptions of such attacks. See, for example, [duncan:etal:2011:statistical, Figure 2.1, pp. 56-58 ].
@@
In our experiments, we use the CEF itself as a labeled database to confirm the accuracy of inferences about {race, ethnicity}. An external attacker would need a similar labeled database to confirm the accuracy of those inferences. It need not be a complete enumeration like the CEF. Small-scale sample surveys or methods like those in [rocher:et:al:2019, ] and [dick:etal:2023, ] could be used. Or the data steward could allow the publication of data about how accurate those inferences would be, as the Census Bureau is doing by releasing the statistics found in this paper. If those inferences are sufficiently more accurate than well-specified baselines, then the confidentiality protections have failed. Such failure is not a {0,1} event. It is a continuum on $[0,1]$. Our methodology calibrates this continuum using the precision of inferences about the features added to the external data from SF1 via the record-linkage attack—specifically, the ratio of confirmed to putative reidentifications. Our vulnerable population is persons who are different from their neighbors on race and ethnicity. We use the census block to define their neighbors. Other vulnerable populations in the 2010 Census include those for whom any characteristic collected on that census differs from their neighbors. See footnote 19 for further elaboration. Our methodology properly distinguishes between vulnerable populations—those where baseline statistical models have low precision and fail to make correct inferences—and nonvulnerable populations—those where baseline statistical models have high precision and generally make correct inferences. We use the results for vulnerable populations to illustrate the confidentiality protection failures in the 2010 Census publications. Specifically, the high reidentification precision rates for vulnerable populations are entirely due to the use of the vulnerable person’s data in the published tables. The most vulnerable persons are population uniques. Delete their data, and the tables that would have contained their responses are completely silent on the risky feature values; that is, the precision of baseline statistical inferences is exactly 0. The vulnerable populations were supposed to be protected by the use of record-level swapping targeting specific population-unique households, but those protections did not recognize how widespread such vulnerable populations were nor how their data could be reconstructed from the ensemble of tables. This confidentiality protection failure can be addressed within a variety of SDL frameworks. We use our reconstruction and reidentification methodology to evaluate three such frameworks: the 2020 Census DAS, which uses a differential privacy framework; enhanced versions of the household swapping framework used for the 1990, 2000 and 2010 Censuses; and the incomplete suppression system used for the 1980 Census.
@@ 5.1. Agreement Match
Algorithm 1 is the basic matching algorithm used generically as part of the agreement, putative, and confirmation matches. Given two databases and a set of common features, Algorithm 1 matches records on the set of features exactly and without replacement.The algorithm iterates over the rows in the left database ($L$) searching, in order, over the rows in the right database ($R$) to look for the first (if any) record that matches on all the selected features. If a matching record is found, the matching records in the left and right databases are both removed and the algorithm continues to the next record in the left database, again looking for a match in the right database. Notice that an essential feature of this matching algorithm is that every record in the left and right databases is at risk for one, and only one, match. It is not possible for a common record type in, say, the right database to be linked to many records in the left database. Failure to enforce this condition results in spurious claims about match rates as, for example, in [ruggles:vanriper:2022, ].

### v2

@@ 4.6. Solution variability implementation details
The differences from the reconstruction in Equation 4 are that equality constraints for tract-level tables become inequality constraints, and the other blocks in tract $t$ are ignored when computing a reconstruction for block $b^{\prime}$. Specifically, only optimization variables for block $b^{\prime}$ are used when forming the tract-level constraints, hence the left-hand side of the tract-level constraint is an underestimate of the counts in the tract-level tables.
@@
Note that the tract-level tables in the solvar Equation 10 appear in inequality constraints. Variables for records in other blocks $b\neq b^{\prime}$ within the same tract $t$ do not appear in this optimization problem for block $b^{\prime}$.
@@
We note that the IP for the initial solution for $\text{rHDF}^{S0}$ is a strict relaxation of the constraints used for the reconstruction $\text{rHDF}_{b,t}$ solving Equation 4. This implies that the feasible region for $\text{rHDF}^{S0}$ is a superset of the feasible region for $\text{rHDF}_{b,t}$. Similarly, the solvar IP in Equation 10 also searches for $\text{rHDF}^{\prime}$ from a strictly larger solution space. This property implies that $\text{max}(2cumsolvar(\text{rHDF}^{S0},\text{B}_{t}),100)$ upper-bounds the solvar value for tract $t$ that an attacker would get using the stronger tract-level constraints. In other words, $\text{max}(2cumsolvar(\text{rHDF}^{S0}$, $\text{B}_{t})$, 100)100)100 ) upper-bounds the percentage of records by which any two solutions to $\text{rHDF}_{b,t}$ can differ because first, the union $\cup_{b\in\text{B}_{t}}\text{rHDF}^{S0}(b)$ is a solution to a problem with a strictly larger solution space than $\text{rHDF}_{b,t}$, and second, $2\sum_{b\in\text{B}_{t}}solvar(\text{rHDF}^{S0},b)$ upper-bounds the percentage of records by which any two solutions to this problem can disagree.
@@
The problem relaxations we adopted, like the use of inequality rather than equality constraints for tract-level tabulations, probably made our upper bound on solution variability much looser. Additionally, we have derived a bound on the distance between any two solutions, but it would be sufficient for most purposes to derive a bound only on the distance between the initial reconstructed solution and an alternative solution, which suggests our solution variability values will often over-state the true solution variability by at least a factor of two. Hence, strengthened (and even more concerning) solution variability bounds may be produced in the future. However, since summing the block-level solution to Equation 10 and multiplying by two is also an upper bound on the exact solution variability for the tract-level problem, the large number of blocks with zero block-level solution variability will still contribute zero to the tract-level solution variability in the $agebin$ schema. Thus, the variability in these blocks is already upper-bounded by zero and cannot get tighter. Tracts with zero tract-level solution variability found in solving the optimization problem IP 10 must also have zero solution variability in any stronger tract-level IP.
@@
We found that solution variability could be readily computed using the optimization in Equation 10 for all 6,207,027 blocks with positive population in the 2010 Census tabulations. This demonstrates that attackers have a public, computationally feasible method for independently identifying blocks for which aggregation into tables introduces no additional uncertainty about the underlying microdata beyond the SDL measures that were applied at the record level to generate the HDF. No access to confidential data is required to perform these solution variability calculations. Furthermore, population uniques found in blocks with zero solution variability are provably population uniques in the HDF, again without using any confidential data. And population uniques using the $agebin$ feature are provably population uniques on the exact age feature. Finally, incorporating additional SF1 tables to the ones shown in Table 2 can only decrease solution variability. Since we used only 18 of the 177 census block-level tables and 16 of the 84 census tract-level tables, our solutions to Equation 10 are also upper bounds on the solution variability of reconstructions that use additional tables.
@@ 5. Matching and Reidentification Methodology
For either $\text{rHDF}_{b,t}$ or $\text{rHDF}_{b}$, our first-order assessment of the quality of the reconstructed microdata matches these reconstructions directly to both the HDF and CEF. This is not a record-linkage assessment. It compares the reconstructed microdata directly to the source microdata for the contingency tables in SF1, namely HDF, and to the source microdata for HDF, namely CEF. We label this step the agreement match because it provides measures of agreement in record-level feature values between $\text{rHDF}_{b,t}$ or $\text{rHDF}_{b}$ and the confidential HDF and CEF. After the agreement match, we do record-linkage based reidentification matches to see how accurately an attacker can infer race and ethnicity from the reconstructed microdata. Thus, we link $\text{rHDF}_{b,t}$ and $\text{rHDF}_{b}$ separately to both commercial data and a specially constructed extract from the CEF called $\text{CEF}_{atkr}$ that includes linking variables (quasi-identifiers) from the schema in Equation 2 and the person identifier pik but no other variables—specifically, not race or ethnicity. These feature sets are shown in the rows COMRCL and $\text{CEF}_{atkr}$ on Table 1.
@@
By record linkage of the quasi-identifiers {block, sex, (age or agebin)} in reconstructed microdata to attacker databases that include names (feature pik), we create putative reidentifications. To enhance the attacker’s database, we attach the data on {race, ethnicity} found on $\text{rHDF}_{b,t}$ or $\text{rHDF}_{b}$ to the {pik, block, sex, (age or agebin)} information in the attacker’s database—either COMRCL or $\text{CEF}_{atkr}$. The feature sets for the records in these putative reidentification databases are shown in the rows labeled Putative $\text{rHDF}_{b,t}$ and Putative $\text{rHDF}_{b}$ in Table 1. Finally, to evaluate the accuracy of the putative reidentifications, and classify confirmed reidentifications, we match putative reidentifications to the full CEF, linking on {pik, block, sex, (age or agebin)} and comparing the {race, ethnicity} inferred from the reconstructed microdata (attached to a putatively reidentified person) to that person’s actual census responses in the CEF. We label the reidentification confirmed, when {pik, block, sex, (age or agebin), race, ethnicity} all match in either the schema in Equation 2 for the feature age or the schema in Table 5 for the feature agebin. The feature sets for these confirmed reidentifications are shown in the rows labeled Confirmed $\text{rHDF}_{b,t}$ and Confirmed $\text{rHDF}_{b}$ in Table 1. Finally, we use the ratio of confirmed to putative reidentifications as the measure of precision or accuracy for the attack.
@@
To make the meaning of this exercise as clear as possible, we state the attacker model we are simulating concisely here. The attacker is an entity external to the Census Bureau with no access to the confidential data contained in HDF or CEF. The attacker has access to all published 2010 Census data—every table in SF1, in particular. The attacker selects a subset (possibly the universe) of these tables and performs record-level reconstruction, possibly using the algorithms in this paper. The attacker also has access to an external database that contains name and address (or some other personal identifying information sufficient to tag a unique person like Social Security Number) and quasi-identifiers—features that match some of the features in SF1—specifically, census block (geocoded from address), sex, and age. Notice that we have explicitly used the same feature set definitions for the attacker’s database as we used in our reconstruction. This is an essential characteristic of a record-linkage attack—the attacker knows enough about the definitions of the feature set for SF1 to construct quasi-identifiers with the same schemas. The attacker matches the external data to the reconstructed record-level data from SF1 using the matching variables (quasi-identifiers) {block, sex, age} and adds the other variables reconstructed from SF1 to the external database. These added variables can include any feature tabulated in SF1, including those tabulated for housing units and householders because those data can be reconstructed as part of the record for “person 1” or “householder,” the individual who completed the census responses for the persons living in the housing unit. In our experiments the extra variables are {race, ethnicity}.
Implementing additional features in the reconstructed data is outside the scope of this paper; however, it is straightforward to add the feature “relationship to householder.” We sketch the details here. See the documentation for Table P29 Household Type by Relationship (sf1:2010:tech, pp. 6-46–6-47). Notice that the universe is “Total Population,” which implies that the marginal total in each of these block-level tables is the block population in Table P1. Create binary variables for the 20 mutually exclusive and exhaustive relationships defined in Table P29: family male householder, family female householder, spouse, biological child, adopted child, stepchild, grandchild, brother or sister, parent, parent-in-law, son-in-law or daughter-in-law, other relatives, nonrelatives, nonfamily male householder living alone, nonfamily male householder not living alone, nonfamily female householder living alone, nonfamily female householder not living alone, nonfamily nonrelatives, group quarters institutionalized person, and group quarters noninstitutionalized person. Add the statistics from Table P29 to the set in Table 2. Add the constraints implied by these tables to the IP in Equations 4 and 10. To further reduce solution variability, if any, in the relationship variables and {race, ethnicity}, add Tables P29A–P29I Household Type by Relationship for major {race, ethnicity} categories. To further reduce the solution variability, if any, in the household population universe, add Table P30. To further reduce the solution variability, if any, for children add Tables P31 and P32. To further reduce the solution variability, if any, for persons ages 65 and older, add Table P34. It is not necessary to add the full variable “relationship to householder” to further reduce solution variability in the features {race, ethnicity}. Using only “lives in household” (all relationships except the last two) and “lives in group quarters” (the last two relationships), use Tables P16A–P16I Population in Households by Age for major {race, ethnicity} categories. Once the relationship “householder” has been reconstructed (all categories that contain the word householder), solution variability for the householder’s {race, ethnicity} can be further reduced by adding Tables P18A–P18I Household Type for major {race, ethnicity} categories. Many tract-level SF1 tables also contain information that can further reduce remaining solution variability.
@@
At this point the attacker is using the information from what we call putative reidentifications as record-level variables on the external database. The attacker may not care about the quality of these putative reidentifications; however, because the solution variability in the schema of our attack is zero for 70% of all census blocks using just the 34 tables in Table 2, the attacker may be satisfied with the 97 million persons for whom they know with certainty they have exactly the information on the confidential internal database. Alternatively, the attacker can keep adding tables from SF1 until solution variability is zero for as large a portion of the target population as desired. For all records for which solution variability is 0, the attacker knows with certainty that the data added from SF1 for those persons match the confidential source records exactly. That is, the attacker knows that they have an image of the confidential Hundred-percent Detail File (HDF) that is completely accurate in the schema of the feature set used for SF1 and the attack. Traditional SDL did not anticipate the possibility of reconstructing accurate microdata from published tables. That is the new feature of our experiments. However, the reidentification attacks we simulate using the record-level reconstructed data exactly match the textbook descriptions of such attacks. See, for example, duncan:etal:2011:statistical.
@@
In our experiments, we use the CEF itself as a labeled database to confirm the accuracy of inferences about {race, ethnicity}. An external attacker would need a similar labeled database to confirm the accuracy of those inferences. It need not be a complete enumeration like the CEF. Small-scale sample surveys or methods like those in rocher:et:al:2019 and dick:etal:2023 could be used. Or the data steward could allow the publication of data about how accurate those inferences would be, as the Census Bureau is doing by releasing the statistics found in this paper. If those inferences are sufficiently more accurate than well-specified baselines, then the confidentiality protections have failed. Such failure is not a {0,1} event. It is a continuum on $[0,1]$. Our methodology calibrates this continuum using the precision of inferences about the features added to the external data from SF1 via the record-linkage attack—specifically, the ratio of confirmed to putative reidentifications. Our vulnerable population is persons who are different from their neighbors on race and ethnicity. We use the census block to define their neighbors. Other vulnerable populations in the 2010 Census include those for whom any characteristic collected on that census differs from their neighbors.
Our methodology properly distinguishes between vulnerable populations—those where baseline statistical models have low precision and fail to make correct inferences—and nonvulnerable populations—those where baseline statistical models have high precision and generally make correct inferences. We use the results for vulnerable populations to illustrate the confidentiality protection failures in the 2010 Census publications. Specifically, the high reidentification precision rates for vulnerable populations are entirely due to the use of the vulnerable person’s data in the published tables. The most vulnerable persons are population uniques. Delete their data, and the tables that would have contained their responses are completely silent on the risky feature values; that is, the precision of baseline statistical inferences is exactly 0. The vulnerable populations were supposed to be protected by the use of record-level swapping targeting specific population-unique households, but those protections did not recognize how widespread such vulnerable populations were nor how their data could be reconstructed from the ensemble of tables. This confidentiality protection failure can be addressed within a variety of SDL frameworks. We use our reconstruction and reidentification methodology to evaluate three such frameworks: the 2020 Census DAS (Disclosure Avoidance System), which uses a differential privacy framework to generate noisy estimates of various census tabulations, then uses mathematical optimization to find microdata that closely approximates these noisy estimates; enhanced versions of the household swapping framework used for the 1990, 2000 and 2010 Censuses; and the incomplete suppression system used for the 1980 Census.
@@ 5.1. Agreement match
Algorithm 1 is the basic matching algorithm used generically as part of the agreement, putative, and confirmation matches. Given two databases and a set of common features, Algorithm 1 matches records on the set of features exactly and without replacement.The algorithm iterates over the rows in the left database ($L$) searching, in order, over the rows in the right database ($R$) to look for the first (if any) record that matches on all the selected features. If a matching record is found, the matching records in the left and right databases are both removed and the algorithm continues to the next record in the left database, again looking for a match in the right database. Notice that an essential feature of this matching algorithm is that every record in the left and right databases is at risk for one, and only one, match. It is not possible for a common record type in, say, the right database to be linked to many records in the left database. Failure to enforce this condition results in spurious claims about match rates as, for example, in ruggles:vanriper:2022.